calculate electric field through the spherical surface of diameter 50cm if the magnitude of the electric field is 1920 N/C and makes an angle of 30degree with the unit vectors


1
Expert's answer
2021-08-05T18:06:40-0400

"\Phi=EA\cos\theta=E\frac{\pi d^2}4\cos \theta=326~\frac{N\cdot m^2}C."